Problem 8: (i) Use the Principle of Mathematical Induction to prove that 2n+1(-1) + 1 1 – 2 + 22 – 23 + ... + (-1)22 = for
Answer #1

problem-8: to prove O ose princple of mathematical nsuchion that 1-272²-23 + .. + (420 = mtl (+1+1 for all positive integersk+ 1 + 3x3 - 2kH (-10k + +(2+1) (kt) kto 3 = 2 kt cook to cesta kit2 + (-1, ktl , ktl +1 = cookt, kt2 + 2 kt (ko - rok 26th +(by ) BUT CH3 = x + 3x² + 3 kt 1 > (2+3) + 3k² takt = 4x² + 3x + 4 7 2k² +2Kity = (K² + 2k to) +3. = (kH)2 +3. = (KiB > (k+1,